OpenStudy (anonymous):

What equation has an x-intercept of 12 and a y-intercept of 6. A. 2x+3y=18 B. -2x+3y=18 C. x+2y=12 D. 2x+2y=4

jimthompson5910 (jim_thompson5910):

if an equation has an x-intercept of 12, then the point (12,0) lies on the line

jimthompson5910 (jim_thompson5910):

so check equation A 2x+3y=18 2(12)+3(0)=18 ... plug in x = 12, y = 0 24 + 0 = 18 24 = 18 ... which is false, so (12,0) does NOT lie on the line 2x+3y = 18

jimthompson5910 (jim_thompson5910):

This shows us that 2x+3y = 18 does NOT have an x-intercept of (12,0)

OpenStudy (anonymous):

x intercept is what you get when you plug in y = 0, and y intercept is what you get when you plug in x = 0

OpenStudy (anonymous):

So plug those in until you find which equation has x = 12 and y = 6.

OpenStudy (anonymous):

okay thanks :)